### v1

@@ 5.2.2 SfM reconstruction results evaluation
As for SuperGlue, its outdoor model demonstrates excellent performance. Even though it might occasionally fail to match some image pairs, it is generally effective at aligning all images and achieves high Feature values, indicating strong performance under the specified conditions. Conversely, its indoor model performs well on Coral-UE5, but underperforms on Coral-2018, with fewer than half of the images successfully aligned. While a limited number of mismatches can be filtered out during the reconstruction process, an excessive number of mismatches poses significant problems. In contrast, LightGlue performs exceptionally well when combined with SIFT or ALIKED, showcasing excellent results. However, its performance is less effective when paired with SuperPoint or DISK. This outcome is largely consistent with the previously obtained image matching results. The results of LoFTR-based DF-SfM are quite distinct, with a high number of Points and a low Track value. This discrepancy can be attributed to the use of detector-free methods, which establish correspondences between images without relying on explicit keypoints. Consequently, the opportunity for repeated observations of the same point is limited. Moreover, its iterative refinement process contributes to the low reprojection errors observed. While it achieves relatively high precision on Coral-UE5, it encounters significant errors on Coral-UE5 (lite), indicating difficulties with reconstruction and insufficient robustness. The results for VGG-SfM are notably different from those of DF-SfM, with VGG-SfM exhibiting a low Points value and a high Track value. This is because that VGG-SfM utilizes deep 2D point tracking to establish correspondences between images. The performance of VGG-SfM on simulation datasets is comparable to that of standard SfM, indicating its viability. However, due to its end-to-end design, VGG-SfM incurs substantial memory overhead, which restricts its use for larger-scale reconstructions. In summary, while SIFT, as a widely-used method, demonstrates solid performance, the learning-based ALIKED generally surpass it, and they are both effective choices. AdaLAM and LightGlue show strong applicability and are also reliable options, though their effectiveness may vary depending on the chosen feature extraction method.
@@ 5.3 Dense surface reconstruction results comparison
Based on the accurate camera poses, dense reconstruction techniques can be employed to capture the intricate geometric structures of coral reefs. This section presents a comparative evaluation of the four categories of dense reconstruction methods discussed in Section 4, focusing on their reconstruction fidelity, accuracy, and efficiency. We implement representative approaches from each category of methods. For the first category, traditional MVS methods, we use the dense reconstruction functionality of COLMAP (Schonberger and Frahm (2016)), which employs a depth map-based MVS algorithm (Schönberger et al. (2016)). In the second category, deep learning-based MVS, we employ Vis-MVSNet (Zhang et al. (2023)) and MVSFormer++ (Cao et al. (2024)). For the third category, NeRF-based methods, we apply Instant-NGP (Müller et al. (2022)), Nerfacto (Tancik et al. (2023)), and Neuralangelo (Li et al. (2023)). Our preliminary experiments indicate that NeuS2 (Wang et al. (2023b)) and BakedSDF (Yariv et al. (2023)) are unable to reconstruct coral reef scenes and are therefore excluded from further comparison. Finally, for the fourth category, GS-based methods, we apply SuGaR (Guédon and Lepetit (2024)), 2D GS (Huang et al. (2024a)), and GOF (Yu et al. (2024)).
It is important to note that different methods produce outputs in various formats (refer to Table 2 in Section 4). For Instant-NGP and Nerfacto, point clouds can be exported using Nerfstudio. The resulting point clouds can then be used with Screened Poisson Surface Reconstruction to generate mesh models (Kazhdan and Hoppe (2013b)). For MVSFormer++, there is a pre-trained model trained on DTU (Aanæs et al. (2016)). as well as a model fine-tuned on Tanks-and-Temples (Knapitsch et al. (2017)), the latter of which performs better overall in our pre-experiments and is therefore used for subsequent comparative analysis. While NeRF-based and GS-based methods have shown promise in dense surface reconstruction, their direct application to large-scale scenes, such as Coral-2018 and Coral-2019, can be challenging. Reconstruction of these scenes can instead be accomplished through MVS methods or by selecting images of a specific region. As for the experimental environment, all experiments in this section were conducted using an NVIDIA GeForce RTX 3090 GPU, and all image data were scaled to 1600 pixels to facilitate processing. COLMAP is implemented in C++, while the other algorithms are implemented in Python. Specifically, Instant-NGP and Nerfacto were executed using the Nerfstudio framework.
@@ 5.3.1 Reconstruction quality evaluation
The direct reconstruction results of various methods on the synthetic image datasets are presented in Figure 10, which includes Coral-UE4, Coral-UE5, and their lite versions. For the Coral-UE4 and Coral-UE5 datasets, the reconstruction results are comparable to those shown in Figure 9. MVSFormer++ produces the most dense point clouds, while Instant-NGP generates sparser point clouds with substantial noise. Both Neuralangelo and 2D GS generate relatively smooth surfaces, whereas GOF exhibits the most superior performance, yielding the most refined and accurate mesh models. In the Coral-UE4 (lite) and Coral-UE5 (lite) datasets, which contain only a quarter of the images compared to the original datasets, the reconstruction results from all solutions exhibit varying degrees of degradation, with noticeable differences in performance. Specifically, COLMAP produces sparser point clouds, with many areas missing points. Although the point clouds reconstructed by Vis-MVSNet and MVSFormer++ are also less complete than those obtained using the full datasets, they are still able to effectively recover the intricate geometric structures of the coral reefs to a reasonable extent. NeRF-based methods exhibit significant performance degradation under limited data conditions. Instant-NGP produces point clouds with a marked increase in noise, impairing the accurate representation of scene geometry. Nerfacto struggles to generate usable point cloud, and Neuralangelo encountered severe errors with the Coral-UE4 (lite) dataset, failing to reconstruct the coral reefs. These issues highlight that NeRF-based approaches depend heavily on datasets with high overlap. When data is sparse, their performance deteriorates sharply, likely due to insufficient data for effective model fitting. In contrast, GS-based methods demonstrate robust performance even with reduced data. SuGaR, 2D GS, and GOF successfully reconstructed the coral reef with high accuracy and completeness. Notably, GOF effectively captured fine coral structures, suggesting that GS-based methods are reliable alternatives when data is limited.
@@
To quantitatively assess the reconstruction accuracy of various methods, we align the 3D points obtained from each method with the ground-truth model from the simulation environments using the Iterative Closest Point (ICP) algorithm. Following the previous work (Schops et al. (2017)), the reconstruction results are evaluated using two metrics: accuracy and completeness. Both measures are evaluated over a range of distance thresholds from 1 to 100. Accuracy indicates how closely the reconstructed points align with the ground-true surface, while completeness measures how much of the actual scene is captured. High accuracy ensures that details are faithfully represented, and high completeness ensures that model covers all important features — both are essential for a reliable 3D reconstruction. Specifically, in this study, accuracy is defined as the percentage of reconstructed 3D points whose distance to the ground-truth mesh is below a given threshold. This distance is calculated using CloudCompare (v2.13.2) with the "Cloud-to-Mesh Distance" tool, which measures the distance from each point in the point cloud to the nearest triangle on the reference mesh. Completeness, on the other hand, is determined by measuring the distance from each ground-truth point to the nearest reconstructed point and is defined as the proportion of ground-truth points that fall within the specified distance threshold. Moreover, given the importance of both metrics, the $F_{1}$ score can serve as a comprehensive single measure to rank the results. The $F_{1}$ score is defined as the harmonic mean of accuracy (precision) $p$ and completeness (recall) $r$, calculated as $2\cdot(p\cdot r)/(p+r)$. The results of each method under different thresholds are shown in Figure 11 and Figure 12. Additionally, Figure 13 displays radar plots of the accuracy, completeness, and $F_{1}$ score for each method. These plots are based on a distance threshold of 10 for the Coral-UE4 and Coral-UE4 (lite) datasets, and a threshold of 20 for the Coral-UE5 and Coral-UE5 (lite) datasets.
@@ 5.3.2 Efficiency
Dense surface reconstruction is typically one of the most time-consuming steps in 3D reconstruction. We calculate the processing time for each method across different datasets, as shown in Table 5, which corresponds to the results presented in Figures LABEL:9 and LABEL:10. Since methods other than COLMAP, Vis-MVSNet, and MVSFormer++ are not yet suitable for reconstructing large-scale scenes, their processing times for Coral-2018 and Coral-2019 were not included in the comparison. It can be observed that Vis-MVSNet and MVSFormer++ are significantly faster than the other methods, taking only a few minutes to process datasets containing dozens of images. In contrast, NeRF-based and GS-based methods require much longer time due to their optimization process, with runtime dependent on the number of training iterations. In this preliminary comparison, the number of iterations for each method was fixed. Although more iterations generally result in better model fitting, the benefits diminish over time, making it necessary to adjust training duration according to specific needs. The efficiency comparison in this paper is preliminary and may change as research progresses. Traditional MVS- and NeRF-based methods are continuously being improved for efficiency, and future research will also focus on enhancing the efficiency of GS-based methods.
@@
Methods Coral-2018 Coral-2019 Coral-2018 (partial) Coral-UE4 Coral-UE4 (lite) Coral-UE5 Coral-UE5 (lite) COLMAP 15246 9531 1287 1910 457 1858 463 Vis-MVSNet 2063 1102 155 183 77 208 81 MVSFormer++ 1498 697 164 172 59 199 61 Instant-NGP - - 1309 1208 1253 1254 1210 Nerfacto - - 1225 1145 1322 1054 1128 Neuralangelo - - 20372 20251 20544 20689 20003 SuGaR - - 9044 6425 5059 5373 5184 2D GS - - 14062 6764 6017 8048 7480 GOF - - 13790 5889 6193 7422 7589
@@ 6.1 Comparison with commercial 3D reconstruction software
To further assess the applicability of recent cutting-edge solutions and provide references for practical engineering applications, we compared two leading commercial 3D reconstruction software packages — Agisoft Metashape and Bentley ContextCapture. These software packages offer functionalities for camera pose estimation and dense reconstruction, and have been used in coral reef reconstruction studies (Bayley and Mogg (2020); Rossi et al. (2020b); Burns and Delparte (2017); Urbina-Barreto et al. (2021a)). For camera pose estimation , we compare the performance of the "ALIKED feature with AdaLAM matching" solution (referred to as ALIKED+AdaLAM) as a benchmark. In Metashape, the “Align Photos” function is used to achieve this, while in ContextCapture, the “Aerotriangulation” process is employed. For dense reconstruction, we employ MVSFormer++ and GOF as reference methods for extracting dense point clouds and mesh models, respectively. In Metashape, the “Build Dense Cloud” and “Build Mesh” functions are used for creating the dense point cloud and mesh model, whereas in ContextCapture, these tasks are carried out by submitting jobs for the “3D point cloud” and “3D mesh” reconstruction. The parameters used in Metashape and ContextCapture are presented in Table 6. Figure 15 illustrates the reconstruction results obtained from processing the Coral-UE4 (lite) dataset with each of these solutions.
@@
Coral-UE4 (lite) Rate 17/17 17/17 15/17
@@
Coral-UE5 (lite) Rate 18/18 18/18 18/18
@@
Figure 16: The boxplots of dense reconstruction errors of different solutions.
@@ 6.2 Coral reef metric estimation via dense surface reconstruction
To assess the ecosystem services and functions of corals, surface area and volume are critical 3D metrics (Zawada et al. (2019)). Surface area is significant as it indicates where coral biomass is concentrated and where coral interacts with its environment (Johannes and Wiebe (1970)). Volume is essential for evaluating the coral reef’s capacity to support biodiversity (Urbina-Barreto et al. (2021b)). Accurate estimation of these metrics is vital for effective coral reef monitoring and conservation. Dense surface reconstruction is a key technology in this process. To assess the effectiveness of different dense reconstruction methods in estimating coral reef metrics, we use the Coral-UE4 dataset as a case study. We produce mesh models of the scene using the approaches involved in Section 5.3 and calculate their surface area and volume. As the reconstructed coral reef mesh models are typically not closed, the holes are filled to create a watertight model firstly. The relative error is then calculated as the difference between estimated and true values, divided by the true value, indicating the accuracy of the estimate. Due to significant variation in surface complexity among different objects, the coral reefs in Coral-UE4 are divided into seven separate objects for statistical analysis, numbered 1 to 7, as illustrated in Figure 17. The relative errors for surface area estimation are presented in Table 8, and for volume estimations in Table 9. We also compute the Root Mean Square Relative Error (RMSRE). It provides an overall assessment of the accuracy of each method. Figure 18 presents the reconstructed mesh model of Object 1 for qualitative analysis.
Figure 17: The boxplots of dense reconstruction errors of different solutions.

### v2

@@ 3.2.2 SfM reconstruction results evaluation
As for SuperGlue, its outdoor model demonstrates excellent performance. Even though it might occasionally fail to match some image pairs, it is generally effective at aligning all images and achieves high Feature values, indicating strong performance under the specified conditions. Conversely, its indoor model performs well on Coral-UE5, but underperforms on Coral-2018, with fewer than half of the images successfully aligned. While a limited number of mismatches can be filtered out during the reconstruction process, an excessive number of mismatches poses significant problems. In contrast, LightGlue performs exceptionally well when combined with SIFT or ALIKED, showcasing excellent results. However, its performance is less effective when paired with SuperPoint or DISK. This outcome is largely consistent with the previously obtained image matching results. The results of LoFTR-based DF-SfM are quite distinct, with a high number of Points and a low Track value. This discrepancy can be attributed to its detector-free design, which establish correspondences between images without relying on explicit keypoints, thereby limiting the potential for repeated observations of the same 3D point. Moreover, its iterative refinement process contributes to the low reprojection errors observed. Although DF-SfM achieves relatively high precision on the Coral-UE5 dataset, it suffers from significant errors on Coral-UE5 (lite), indicating challenges in reconstruction and limited robustness under more constrained conditions. Furthermore, despite its very low reprojection errors on Coral-2018, it exhibits large ground control point (GCP) errors, suggesting inconsistencies in global alignment. The results for VGG-SfM are notably different from those of DF-SfM, with VGG-SfM exhibiting a low Points value and a high Track value. This is because that VGG-SfM utilizes deep 2D point tracking to establish correspondences between images. The performance of VGG-SfM on simulation datasets is comparable to that of standard SfM, indicating its viability. However, due to its end-to-end design, VGG-SfM incurs substantial memory overhead, which restricts its use for larger-scale reconstructions. In summary, while SIFT, as a widely-used method, demonstrates solid performance, the learning-based ALIKED generally surpass it, and they are both effective choices. AdaLAM and LightGlue show strong applicability and are also reliable options, though their effectiveness may vary depending on the chosen feature extraction method.
@@ 3.3 Dense surface reconstruction results comparison
Based on the accurate camera poses, dense reconstruction techniques can be employed to capture the intricate geometric structures of coral reefs. This section presents a comparative evaluation of the four categories of dense reconstruction methods discussed in Section 2.3, focusing on their reconstruction fidelity, accuracy, and efficiency. We implement representative approaches from each category of methods. For the first category, traditional MVS methods, we use the dense reconstruction functionality of COLMAP (Schonberger and Frahm (2016)), which employs a depth map-based MVS algorithm (Schönberger et al. (2016)). In the second category, deep learning-based MVS, we employ Vis-MVSNet (Zhang et al. (2023)) and MVSFormer++ (Cao et al. (2024)). For the third category, NeRF-based methods, we apply Instant-NGP (Müller et al. (2022)), Nerfacto (Tancik et al. (2023)), and Neuralangelo (Li et al. (2023)). Our preliminary experiments indicate that NeuS2 (Wang et al. (2023b)) and BakedSDF (Yariv et al. (2023)) are unable to reconstruct coral reef scenes and are therefore excluded from further comparison. Finally, for the fourth category, GS-based methods, we apply SuGaR (Guédon and Lepetit (2024)), 2D GS (Huang et al. (2024a)), and GOF (Yu et al. (2024)).
It is important to note that different methods produce outputs in various formats (refer to Table 2 in Section 2.3). For Instant-NGP and Nerfacto, point clouds can be exported using Nerfstudio. The resulting point clouds can then be used with Screened Poisson Surface Reconstruction to generate mesh models (Kazhdan and Hoppe (2013)). For MVSFormer++, there is a pre-trained model trained on DTU (Aanæs et al. (2016)) as well as a model fine-tuned on Tanks-and-Temples (Knapitsch et al. (2017)), the latter of which performs better overall in our pre-experiments and is therefore used for subsequent comparative analysis. While NeRF-based and GS-based methods have shown promise in dense surface reconstruction, their direct application to large-scale scenes, such as Coral-2018 and Coral-2019, can be challenging. Reconstruction of these scenes can instead be accomplished through MVS methods or by selecting images of a specific region. As for the experimental environment, all experiments in this section were conducted using an NVIDIA GeForce RTX 3090 GPU, and all image data are scaled to a maximum dimension of 1600 pixels. COLMAP is implemented in C++, while the other algorithms are implemented in Python. Specifically, Instant-NGP and Nerfacto were executed using the Nerfstudio framework.
@@ 3.3.1 Reconstruction quality evaluation
The direct reconstruction results of various methods on the synthetic image datasets are presented in Figure 11, which includes Coral-UE4, Coral-UE5, and their lite versions. For the Coral-UE4 and Coral-UE5 datasets, the reconstruction results are comparable to those shown in Figure 10. MVSFormer++ produces the most dense point clouds, while Instant-NGP generates sparser point clouds with substantial noise. Both Neuralangelo and 2D GS generate relatively smooth surfaces, whereas GOF exhibits the most superior performance, yielding the most refined and accurate mesh models. In the Coral-UE4 (lite) and Coral-UE5 (lite) datasets, which contain only a quarter of the images compared to the original datasets, the reconstruction results from all solutions exhibit varying degrees of degradation, with noticeable differences in performance. Specifically, COLMAP produces sparser point clouds, with many areas missing points. Although the point clouds reconstructed by Vis-MVSNet and MVSFormer++ are also less complete than those obtained using the full datasets, they are still able to effectively recover the intricate geometric structures of the coral reefs to a reasonable extent. NeRF-based methods exhibit significant performance degradation under limited data conditions. Instant-NGP produces point clouds with a marked increase in noise, impairing the accurate representation of scene geometry. Nerfacto struggles to generate usable point clouds, and Neuralangelo encountered severe errors with the Coral-UE4 (lite) dataset, failing to reconstruct the coral reefs. These issues highlight that NeRF-based approaches depend heavily on datasets with high overlap. When data is sparse, their performance deteriorates sharply, likely due to insufficient data for effective model fitting. In contrast, GS-based methods demonstrate robust performance even with reduced data. SuGaR, 2D GS, and GOF successfully reconstructed the coral reef with high accuracy and completeness. Notably, GOF effectively captured fine coral structures, suggesting that GS-based methods are reliable alternatives when data is limited.
@@
To quantitatively assess the reconstruction accuracy of various methods, we align the 3D points obtained from each method with the ground-truth model from the simulation environments using the Iterative Closest Point (ICP) algorithm. Following the previous work (Schops et al. (2017)), the reconstruction results are evaluated using two metrics: accuracy and completeness. Both measures are evaluated over a range of distance thresholds from 1 to 100. Accuracy indicates how closely the reconstructed points align with the ground-true surface, while completeness measures how much of the actual scene is captured. High accuracy ensures that details are faithfully represented, and high completeness ensures that model covers all important features — both are essential for a reliable 3D reconstruction. Specifically, in this study, accuracy is defined as the percentage of reconstructed 3D points whose distance to the ground-truth mesh is below a given threshold. This distance is calculated using CloudCompare (v2.13.2) with the "Cloud-to-Mesh Distance" tool, which measures the distance from each point in the point cloud to the nearest triangle on the reference mesh. Completeness, on the other hand, is determined by measuring the distance from each ground-truth point to the nearest reconstructed point and is defined as the proportion of ground-truth points that fall within the specified distance threshold. Moreover, given the importance of both metrics, the $F_{1}$ score can serve as a comprehensive single measure to rank the results. The $F_{1}$ score is defined as the harmonic mean of accuracy (precision) $p$ and completeness (recall) $r$, calculated as $2\cdot(p\cdot r)/(p+r)$. The results of each method under different thresholds are shown in Figure 12 and Figure 13. Additionally, Figure 14 displays radar plots of the accuracy, completeness, and $F_{1}$ score for each method. These plots are based on a distance threshold of 10 for the Coral-UE4 and Coral-UE4 (lite) datasets, and a threshold of 20 for the Coral-UE5 and Coral-UE5 (lite) datasets. Additionally, the comparative performance of various methodologies is visualized in Figure 14 through radar plots displaying three key evaluation metrics: accuracy, completeness, and $F_{1}$ score. To accommodate the variability in camera-to-object distances across experimental scenarios, different distance thresholds are established: 10 for the Coral-UE4 and Coral-UE4 (lite) datasets, and 20 for the Coral-UE5 and Coral-UE5 (lite) datasets.
@@ 3.3.2 Efficiency
Dense surface reconstruction is typically one of the most time-consuming steps in 3D reconstruction. We calculate the processing time for each method across different datasets, as shown in Table 5. Since methods other than COLMAP, Vis-MVSNet, and MVSFormer++ are not yet suitable for reconstructing large-scale scenes, their processing times for Coral-2018 and Coral-2019 were not included in the comparison. It can be observed that Vis-MVSNet and MVSFormer++ are significantly faster than the other methods, taking only a few minutes to process datasets containing dozens of images. In contrast, NeRF-based and GS-based methods require much longer time due to their optimization process, with runtime dependent on the number of training iterations. In this preliminary comparison, the number of iterations for each method was fixed. Although more iterations generally result in better model fitting, the benefits diminish over time, making it necessary to adjust training duration according to specific needs. The efficiency comparison in this paper is preliminary and may change as research progresses. Traditional MVS- and NeRF-based methods are continuously being improved for efficiency, and future research will also focus on enhancing the efficiency of GS-based methods.
@@
Methods Coral-2018 Coral-2019 Coral-2018 (partial) Coral-UE4 Coral-UE4 (lite) Coral-UE5 Coral-UE5 (lite) COLMAP 13565 9531 1287 1910 457 1858 463 Vis-MVSNet 1817 1102 155 183 77 208 81 MVSFormer++ 1311 697 164 172 59 199 61 Instant-NGP - - 1309 1208 1253 1254 1210 Nerfacto - - 1225 1145 1322 1054 1128 Neuralangelo - - 20372 20251 20544 20689 20003 SuGaR - - 9044 6425 5059 5373 5184 2D GS - - 14062 6764 6017 8048 7480 GOF - - 13790 5889 6193 7422 7589
@@ 3.4 Comparison with commercial 3D reconstruction software
To further assess the applicability of recent cutting-edge solutions and provide references for practical engineering applications, we compared two leading commercial 3D reconstruction software packages — Agisoft Metashape and Bentley ContextCapture. These software packages offer functionalities for camera pose estimation and dense reconstruction, and have been used in coral reef reconstruction studies (Bayley and Mogg (2020); Rossi et al. (2020); Burns and Delparte (2017); Urbina-Barreto et al. (2021)). For camera pose estimation, we compare the performance of the "ALIKED feature with AdaLAM matching" solution (referred to as ALIKED+AdaLAM) as a benchmark. In Metashape, the “Align Photos” function is used to achieve this, while in ContextCapture, the “Aerotriangulation” process is employed. For dense reconstruction, we employ MVSFormer++ and GOF as reference methods for extracting dense point clouds and mesh models, respectively. In Metashape, the “Build Dense Cloud” and “Build Mesh” functions are used for creating the dense point cloud and mesh model, whereas in ContextCapture, these tasks are carried out by submitting jobs for the “3D point cloud” and “3D mesh” reconstruction. The parameters used in Metashape and ContextCapture are presented in Table 6. Figure 16 illustrates the reconstruction results obtained from processing the Coral-UE4 (lite) dataset with each of these solutions.
@@
Coral-UE4 (lite)
Rate 17/17 17/17 15/17
@@
Coral-UE5 (lite)
Rate 18/18 18/18 18/18
@@
Figure 17: Accuracy and completeness of the reconstruction results for each approach at different distance thresholds.
@@ 3.5 Coral reef metric estimation via dense surface reconstruction
To assess the ecosystem services and functions of corals, surface area and volume are critical 3D metrics (Zawada et al. (2019)). Surface area is significant as it indicates where coral biomass is concentrated and where coral interacts with its environment (Johannes and Wiebe (1970)). Volume is essential for evaluating the coral reef’s capacity to support biodiversity (Urbina-Barreto et al. (2021)). Accurate estimation of these metrics is vital for effective coral reef monitoring and conservation. Dense surface reconstruction is a key technology in this process. To assess the effectiveness of different dense reconstruction methods in estimating coral reef metrics, we use the Coral-UE4 dataset as a case study. We produce mesh models of the scene using the approaches involved in Section 3.3 and calculate their surface area and volume. As the reconstructed coral reef mesh models are typically not closed, the holes are filled to create a watertight model firstly. The relative error is then calculated as the difference between estimated and true values, divided by the true value, indicating the accuracy of the estimate. Due to significant variation in surface complexity among different objects, the coral reefs in Coral-UE4 are divided into seven separate objects for statistical analysis, numbered 1 to 7, as illustrated in Figure 18. The relative errors for surface area estimation are presented in Table 8, and for volume estimations in Table 9. We also compute the Root Mean Square Relative Error (RMSRE). It provides an overall assessment of the accuracy of each method. Figure 19 presents the reconstructed mesh model of Object 1 for qualitative analysis.
Figure 18: The IDs of objects in the Coral-UE4.
@@
4 Discussions
4.1 Accuracy and robustness of camera pose estimation
The primary objective in camera pose estimation is to compute accurate and robust camera parameters for all images within a dataset. Our experiments (Section 3.2) show that not all cutting edge matching methods guarantee improved performance across all metrics, but several recent techniques demonstrate clear benefits in specific scenarios. Overall, certain emerging techniques offer compelling advantages.
For example, when using SIFT as the feature detector, LightGlue achieves more accurate overall reconstructions, notably on the Coral UE5 (lite) dataset. Although LightGlue sometimes produces a slightly higher reprojection error compared to the traditional ratio test, this does not directly correlate with object space accuracy. LightGlue’s contribution led to a more complete and globally consistent structure, making it a superior choice for the overall reconstruction pipeline. Similarly, ALIKED proves to be an effective feature detector, offering a level of scale invariance that many deep learning based descriptors struggle to match. Its robustness greatly increases the success rate of image alignment, positioning it as a strong alternative.
To improve the success rate and quality of the subsequent dense reconstruction phase in the overall workflow, it is important to align as many images as possible during the SfM stage. Our findings across different datasets reveal that so long as the images are successfully aligned, the final accuracy tends to converge to a comparable level. For instance, multiple pipelines achieve millimeter-level error of the check point on Coral-2018 and similar pose errors on Coral-UE5. These suggest that the current bottleneck lies not in the precision of feature localization but in the robustness of feature matching. Specifically, the ability to establish reliable correspondences, especially under challenging conditions. The comparison with commercial software in Section 3.4 supports this view, where ContextCapture’s failure to align all images on the Coral-UE4 dataset led to significantly larger errors.
The promising performance of methods like SuperGlue and LightGlue underscores the potential of deep learning-based matching techniques. These methods leverage broader image context to generate more reliable correspondences than traditional approaches. This success suggests that further exploration of these avenues is valuable. Furthermore, exploring the integration of prior knowledge, like known image sequences or approximate geolocation, offers another potential pathway to bolster alignment robustness, especially within large or complex scenes.
4.2 Dense surface reconstruction with enhanced quality and efficiency
Regarding dense surface reconstruction, our experiments highlight the transformative impact of recent advances like NeRF and 3D GS. These techniques have spurred the development of novel methods capable of achieving high-quality reconstructions, capturing intricate geometric details and realistic appearances. Concurrently, deep learning-based MVS methods, which integrate deep learning components into typical MVS pipelines, deliver strong performance with favorable runtimes. As technology evolves, we anticipate further breakthroughs in both approaches.
In practical applications, particularly for large-scale surveys common in environmental monitoring (e.g., coral reef mapping), the trade-off between reconstruction quality and computational efficiency is a critical consideration. Based on the results presented in Section 3.3, deep learning-based MVS methods, particularly MVSFormer++, achieve an excellent balance. They offer a compelling general-purpose solution, delivering high-quality results with reasonable computational costs. While methods like GOF can achieve outstanding accuracy and completeness, their significant computational cost may render them less practical for extensive areas.
However, for coral reef 3D reconstruction specifically, a single method may not be optimal for the entire scene. Reef environments often contain diverse structures: complex coral colonies requiring high fidelity, alongside simpler areas like rock formations or sandy patches where extreme detail is less critical. This heterogeneity suggests a potential hybrid strategy. For example, one could employ an efficient method like MVSFormer++ for broad-scale reconstruction of the entire survey area. Subsequently, specific regions of interest containing intricate or scientifically significant coral structures could be selectively reconstructed using a high-fidelity but more computationally intensive method like GOF. The resulting high-detail local models could then be integrated into global reconstruction. This offers a practical pathway to balance overall quality and efficiency. Developing robust methodologies for planning such hybrid reconstructions, segmenting regions for differential processing, and seamlessly fusing the results from different methods represents a valuable direction for future research, potentially leading to optimized workflows tailored for complex, large-scale 3D modeling tasks.